OpenStudy (anonymous):

what is 3x-5y=9,x=-1,0,1

OpenStudy (anonymous):

Substitute each x value into the given equation: 3(-1)-5y=9 solve for y -3-5y=9 -5y=12 y=-12/5 Do that process for the remaining two x values 0 & 1

OpenStudy (anonymous):

Can you take it from here

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

which part

OpenStudy (anonymous):

What do i do with the Y????

OpenStudy (anonymous):

how do u get -5y = 12?

OpenStudy (anonymous):

3(-1)-5y=9 -3-5y=9 -3+3-5y =9+3 added 3 to each side to get -5y by itself 0-5y=12 simplified -3+3=0 -5y=12 simplified 0-5y=5y

OpenStudy (anonymous):

-5y=12 -5y/-5=12/-5 getting y by itself y=-12/5 so when x is -1, y is -12/5 (-1,-12/5)

OpenStudy (anonymous):

Now let's figure out 3x-5y=9 when x=0 3(0)-5y=9 substituted in where I saw the x 0-5y=9 simplified -5y=9 simplified -5y/-5=9/-5 divided by -5 to get y by itself y=-9/5 when x is 0, y is -9/5 (0,-9/5)

OpenStudy (anonymous):

Now, you solve the equation when x=1 follow the same process

OpenStudy (anonymous):

ok thxxxx

OpenStudy (anonymous):

do you understand

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

OK...x=1 substitute into the equation 3x-5y=9 put 1 in where you see an x 3(1)-5y=9 do you understand this part

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

OK...simplify this...multiply out the 3(1) 3(1)-5y=9 3-5y=9 Understand this part

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

whats the next part???

OpenStudy (anonymous):

after 3(0) part

OpenStudy (anonymous):

the 3(0) means to multiply 3 and 0, which gives us 0

OpenStudy (anonymous):

still there

OpenStudy (anonymous):

Later

OpenStudy (texaschic101):

you got a lot of patience DSS :) Cheers to you

OpenStudy (anonymous):

thanks